OpenStudy (anonymous):

The weight (W) of an object is the product of its mass (m) in kilograms and g, the acceleration due to gravity in meters/second2. If the acceleration due to gravity is 9.8 meters/second2, find the corresponding weights for these masses. mass in kilograms = {22.1, 33.5, 41.3, 59.2, 78} The weights (in newtons) for the given masses are

OpenStudy (astrophysics):

\[W = m \times g\] so what's the problem herre?

OpenStudy (astrophysics):

They have given you the mass, and the gravity which is 9.8 m/s^2

OpenStudy (astrophysics):

Paying attention to the units, you can see Newton's is the product of kg multiplied by m/s^2. \[W = m \times g = kg \times \frac{ m }{ s^2 } = N\]

OpenStudy (anonymous):

so whats the answer?

OpenStudy (astrophysics):

We don't give direct answers here, I'm trying to help you understand the concept so you can do it on your own.

OpenStudy (astrophysics):

You want weight in newtons so just multiply each of your masses by gravity

OpenStudy (anonymous):

ok so what do I do

OpenStudy (astrophysics):

\[W_1 = 22.1 \times 9.8 = ...\] etc for all your weights

OpenStudy (astrophysics):

where the subscript 1 represents your first weight using first mass

OpenStudy (anonymous):

oh ok